### v1

@@ Abstract
The Agatston score, which is the sum of the calcification in the four main coronary arteries, has been widely used in the diagnosis of coronary artery disease (CAD). However, many studies have emphasized the importance of the vessel-specific Agatston score, as calcification in a specific vessel is significantly correlated with the occurrence of coronary heart disease (CHD). In this paper, we propose the Residual-block Inspired Coordinate Attention U-Net (RICAU-Net), which incorporates coordinate attention in two distinct manners and a customized combo loss function for lesion-specific coronary artery calcium (CAC) segmentation. This approach aims to tackle the high class-imbalance issue associated with small and sparse lesions, particularly for CAC in the left main coronary artery (LM) which is generally small and the scarcest in the dataset due to its anatomical structure. The proposed method was compared with six different methods using Dice score, precision, and recall. Our approach achieved the highest per-lesion Dice scores for all four lesions, especially for CAC in LM compared to other methods. The ablation studies demonstrated the significance of positional information from the coordinate attention and the customized loss function in segmenting small and sparse lesions with a high class-imbalance problem.
Keywords:
Coordinate Attention Small lesion Coronary artery calcium
@@ 1 Introduction
Cardiovascular disease (CVD) has constantly been one of the leading causes of death globally for decades [33]. Among the components of CVD, coronary artery disease (CAD) is highly vital leading to sudden death as the coronary arteries are narrowed and blocked due to the coronary artery calcification. Currently, coronary artery calcium (CAC) can be non-invasively detected by using non-contrast cardiac computed tomography (CT) [33, 19]. By analyzing the collected CT images, the CAC score, known as the Agatston score [1], is calculated to quantify the severity of CAD [9, 41]. The Agatston score is the sum of calcification in the four main coronary arteries: the left main coronary artery (LM), left anterior descending artery (LAD), left circumflex artery (LCX), and right coronary artery (RCA). As a result, it is markedly correlated with the occurrence of CVD while recognized as a strong CVD prediction index [39, 25, 12, 32]. With the advancement of artificial intelligence in the medical field, deep learning algorithms have shown its capability of locating the calcium in the coronary arteries and automating the calculation of the CAC score with high precision [24, 15, 5, 21, 31]. Although the total Agatston score has been drawing its attention as a strong CVD prediction index for decades, many studies [15, 31, 35, 34, 43] emphasized the importance of the lesion-specific CAC scores. The authors in [36, 37, 20] demonstrated that the calcification in the LM increased the CVD death and all-cause mortality. Besides, the authors in [3] concluded that the presence of CAC in RCA was significantly correlated with coronary heart disease (CHD) and CVD. Therefore, the importance of the per-lesion Agatston score is as significant as the conventional Agatston score in predicting CHD and CVD. To calculate the per-lesion Agatston score, the lesion-specific CAC segmentation is a mandatory process. However, there are two limitations when facilitating the lesion-specific segmentation of CAC. The first factor to consider is the size and sparsity of the lesions. CAC is located in the coronary arteries; therefore, the size of CAC is relatively smaller than that of other organs such as the lung, kidney, or liver. The second issue is the high class-imbalance problem in the datasets. Several studies [2, 42, 4] have shown that CAC in LM is the least common lesion among the four main coronary arteries. This is because LM is the shortest among coronary arteries, especially in patients with coronary atherosclerosis [7, 8]. Therefore, the number of CT slices where CAC in LM appeared is the smallest, and the size of the LM lesion is also small. Among the CAC segmentation studies [24, 15, 5, 21, 31] that utilized deep learning algorithms, three studies [15, 21, 31] focused on lesion-specific segmentations. However, [21, 31] combined CAC in LM and LAD as one lesion due to the challenge of distinguishing CAC in LM from CAC in LAD.
@@
Herein, we propose the Residual-block Inspired Coordinate Attention U-Net (RICAU-Net), an end-to-end deep learning algorithm for lesion-specific CAC segmentation using non-contrast cardiac CT scans. We also propose a customized combo loss function to tackle the segmentation of small and sparse lesions with a high class-imbalance problem. In the reporting deep learning algorithm, coordinate attention (CA) [16] was implemented as RICA blocks in the encoder and as CA modules in the decoder of the proposed model. This implementation aimed to highlight the positional information of CAC as the locations of CAC were closely related to their surrounding vessels in CT images, which is substantial for the lesion-specific segmentation task.
Our main contributions in this paper are summarized as below:
  1. 1.
    We propose the Residual-block Inspired Coordinate Attention U-Net and the customized combo loss function. RICAU-Net focuses on the positional information of small and sparse CAC lesions by utilizing the RICA blocks and CA modules. The proposed loss function is utilized to tackle the segmentation of small and sparse lesions with a high class-imbalance issue.
  2. 2.
    We demonstrated the effectiveness of the RICA blocks, CA modules, and the proposed loss function in lesion-specific CAC segmentation, particularly for CAC in LM which was small and the scarcest in the dataset.
2 Related Works
2.1 Medical Image Segmentation Models
U-Net is one of the most ubiquitous neural network architectures for semantic segmentation, originally developed for biomedical image segmentation [29]. U-Net is a U-shaped encoder-decoder type fully convolutional network (FCN) that can accept input images of any size. The encoder and the decoder are connected with skip connections that concatenate the feature maps from the encoder to the decoder to aggregate contextual information and local information from different layers. Due to its outstanding performance in image segmentation tasks, many variants were developed for and showed excellent results in segmentation of medical images, such as SA U-Net [10], Nested U-Net [44], ResUNet++ [18], Attention U-Net [26], and classical network backboned U-Net such as EfficientNet U-Net, ResNet U-Net, VGG U-Net.
2.2 Coordinate Attention
Attention mechanisms have been actively researched in the field of computer vision over the past years because they direct the models to focus on specific areas in the image. The attention mechanisms are categorized into three groups: channel attention, spatial attention, and combined attention [11]. Channel attention focuses on the global information of an input image by utilizing the channel-connection among features maps, whereas spatial attention focuses on the regional information of an input image by utilizing the spatial-connection among features maps [40]. Combined attention is the fusion of channel and spatial attention. The integration of these mechanisms into deep learning models has been proven to enhance the performance of the models. SENet (channel) [17], CBAM (mixed) [40], and Vision Transformers [6] are the representative examples of attention-based models. However, the mechanisms mentioned above did not incorporate positional information from the input image. Thus, Hou et al. [16] designed coordinate attention for mobile networks to aggregate the positional information from the input image into channel attention. The CA module generates direction-aware and position-sensitive feature maps by breaking down channel attention into two 1D feature encoding processes that retain information in both horizontal and vertical directions. These feature maps help the neural network accurately detect and locate regions of interest in the input image.
Figure 1: Overview of the RICAU-Net architecture. RICAU-Net was designed based on the vanilla U-Net [42]. In the encoder, the RICA block utilizes coordinate attention information from the input in each layer (Fig. 2). In the decoder, four CA modules were implemented after each upsampling process.
3 Method
@@ 3.1 Coordinate Attention Module
The locations and types of CAC mainly depend on the surrounding coronary arteries. Therefore, the positional information of CAC and the characteristics of the surrounding vessels can provide crucial information for lesion-specific CAC segmentation. In addition, CAC in LM is typically smaller and scarcer compared to CAC in other coronary arteries due to its anatomical structure [2, 4]. And they are not easily distinguishable when the nearby calcium lesions are from multiple coronary arteries and coexist as a cluster. In such cases, the positional information of CAC becomes more crucial when segmenting the calcium lesions. Hence, the CA module was incorporated into the proposed architecture to emphasize the positional information of CAC and its surrounding arteries from the input images. The CA module is defined as:
@@
Figure 2: Overview of the CA module and the encoder block in RICAU-Net (RICA block). The CA module illustrates the overall process of aggregating positional information from both horizontal and vertical directions of a feature map into channel attention. In the RICA block, the final output is obtained by performing an element-wise addition operation of the output of the two convolutional layers and the output of the skip connection. The skip connection comprises a CA module for direction-aware and position-sensitive feature map generation and a projection shortcut (PJS) module for dimension matching. In contrast, the encoder block in U-Net only consists of two convolutional layers.
@@ 3.2 Residual-block Inspired Coordinate Attention U-Net
The decoder of the RICAU-Net is similar to that of the vanilla U-Net, but it incorporates four CA modules after each upsampling process. In this way, the network is capable of retaining the positional information of the upsampled features. The RICA block was not utilized in the decoder of the RICAU-Net because it could result in the loss of essential positional information when the feature maps from the CA module pass through the PJS module for dimension matching. This decision was also empirical, as the proposed architecture demonstrated better segmentation performance, particularly on CAC in LM, compared to the architecture with the RICA block in the decoder during the initial experiment with the architecture design.
@@ 3.3 Class-Imbalance Problem and Loss Function
3.3.1 Class-Imbalance Problem.
Our non-contrast cardiac CT scan dataset shows a significant class imbalance between the negative samples (background, bone) and the positive samples (CAC in LM, LAD, LCX, RCA). For instance, the training set of 494 scans was sliced into $512\times 512$ pixel resolution of 29,154 images. Among the 29,154 images, the smallest number of CAC was found in LM, appearing in only 390 images (1.3% of the training set), whereas the largest number of CAC was observed in RCA, presenting in 2,144 images (7.4% of the training set). This phenomenon persisted in the validation and test sets as well. Among the validation and test sets, only 1.3% and 0.4% of images contained CAC in LM, respectively. This phenomenon became more significant at the pixel level. In a $512\times 512$ pixel image, one LM lesion (positive sample) could be as small and sparse as 5 pixels $(>1mm^{2})$, whereas the remaining pixels would belong to the negative samples (background or bone).
3.3.2 Loss Function.
Our dataset suffers from two major problems. The first issue is high class imbalance. The second issue is the small and sparse sizes of the target lesions. Focal loss was first introduced by Lin et al. [22] to address the class imbalance problem in the object detection field by focusing more on hard examples while preventing easy negatives from influencing the model during training. Medical datasets often face the issue of high class-imbalance problem due to their inherent nature. Qiao et al. [27] and Qin et al. [28] employed focal loss to tackle the class imbalance problem in x-ray classification tasks, resulting in improved results. Exponential Logarithmic Loss (ELL) was introduced by Wong et al. [38] to address the highly unbalanced object size problem in 3D segmentation. ELL is a type of combo loss that is based on a mixture of Cross-entropy loss and Dice loss. Wong et al. [38] used logarithmic Dice loss to focus more on poorly segmented objects and combined it with weighted Cross-entropy loss. When combining, exponential parameters were applied to both loss functions to further impose the nonlinearities for accurate segmentation of small and large objects.
In this study, to overcome the two issues, we employed a weighted Focal LogDice loss, which combined a weighted Focal loss with an exponential logarithmic Dice loss. For the weighted Focal loss, the weighting factor in the multiclass scenario is set as the inverse class frequency of the number of images per class. The weighting factor of the class $k_{c}$ is defined as:
$\alpha_{k_{c}}=\frac{\sum\nolimits_{c=0}^{k_{c}}{I_{k_{c}}}}{({c_{max}}+1)\ast I_{k_{c}}}$ (3)
where $\alpha_{k_{c}}$ denotes the weighting factor of the class, $k_{c}$ denotes the class $k$ with a class index $c$, where $c=[0,5]$ as our data has six classes, and $I_{k_{c}}$ denotes the number of images per class. The final weighted focal loss is defined as:
$\mathcal{L}_{WFL}=\frac{\sum\nolimits_{n=1}^{N}{\sum\nolimits_{c=0}^{k_{c}}{(-\alpha_{k_{c}}(1-p_{t,k_{c}})^{\gamma}\log{p_{t,k_{c}}})}}}{N}$ (4)
where $\mathcal{L}_{WFL}$ denotes weighted focal loss, $N$ denotes for the number of pixels, $n$ denotes for notation for pixel, $p_{t,{k_{c}}}$ denotes the prediction probability of pixel per class, $\gamma$ denotes the tunable focusing parameter and $\gamma=2$ was used as suggested by Lin et al. [22]. For the exponential logarithmic Dice loss, the loss function is defined as:
$\mathcal{L}_{ELDL}=\left(-\ln\frac{2*\sum\nolimits_{n=1}^{N}{\sum\nolimits_{c=0}^{k_{c}}{p_{true,k_{c}}\ast p_{t,k_{c}}}}}{\sum\nolimits_{n=1}^{N}{\sum\nolimits_{c=0}^{k_{c}}{p_{true,k_{c}}}}+\sum\nolimits_{n=1}^{N}{\sum\nolimits_{c=0}^{k_{c}}{p_{t,k_{c}}}}}\right)^{\gamma_{Dice}}$ (5)
where $\mathcal{L}_{ELDL}$ denotes exponential logarithmic Dice loss, $p_{true,{k_{c}}}$ denotes the ground truth probability of pixel per class, $\gamma_{Dice}$ is the nonlinearity control parameter and $\gamma_{Dice}=0.3$ was used as it yielded the best result for Wong et al. [38]. The final weighted Focal LogDice loss is defined as:
@@
where $\mathcal{L}_{WFLogDL}$ denotes the weighted Focal LogDice loss, $w_{Focal}$ and $w_{Dice}$ denote weights of each loss function, and $w_{Focal}=0.4$ and $w_{Dice}=0.6$ were used in this study.
4 Experiment
4.1 Dataset, Preprocessing, and Implementation
4.1.1 Dataset.
The collection of clinical CT images and informed consent form were reviewed and approved by the institutional review board of SingHealth (IRB 2022-2521). A total of 1,108 non-contrast cardiac CT scans were used in this study, acquired using four different scanners: SIEMENS, GE Medical Systems, PHILIPS, and TOSHIBA. The calcium lesions in the coronary arteries were manually annotated by five clinically trained specialists to create lesion-specific ground truth masks. The ground truth masks contained multiclass information on background, bone, CAC in LM, LAD, LCX, and RCA. The CT scans and their corresponding ground truth masks were divided into training, validation, and test sets. A total of 612 CT scans from SIEMENS, GE Medical Systems, and Philips scanners were used for the training and validation sets. The training set (80%) comprised 494 scans (484 subjects with CAC and 10 subjects without CAC), and the validation set (20%) comprised 118 scans (112 subjects with CAC and 6 subjects without CAC). 496 CT scans from TOSHIBA scanner were used for the test set, comprising 278 subjects with CAC and 218 subjects without CAC. All the CT scans had the same height and width dimensions ($512\times 512$ pixel sizes).
4.1.2 Preprocessing.
All the CT scans were sliced into 2D images with a resolution of $512\times 512$ pixels. The raw Hounsfield Units (HU) of each slice were clipped to a range of [-150, 230], which represents the clinical CT windowing of the heart (window width: 380, window level: 40), followed by normalization to a range of floating-point numbers [0, 1]. In addition, random rotations of 5 and 10 degrees clockwise and counterclockwise, random center crops of $300\times 300$ and $400\times 400$ pixel resolutions, as well as the addition of Gaussian blur, Gaussian noise, and salt and pepper noise, were employed as on-the-fly data augmentation methods to augment the training data. When a random center crop was applied, the image was padded to a $512\times 512$ pixel resolution.
4.2 Implementation Details and Evaluation Metrics
4.2.1 Implementation Details.
@@ 4.2 Implementation Details and Evaluation Metrics
4.2.2 Evaluation Metrics.
Dice score, precision, and recall were selected as the evaluation metrics to access the per-lesion performance of the models at the pixel level. Dice score is one of the most commonly used evaluation metrics in semantic segmentation tasks because it measures the similarity between the ground truth and the model’s prediction. The selected evaluation metrics can be easily calculated by the combinations of true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN). The Dice score is calculated as 2TP/(2TP+FP+FN), precision as TP/(TP+FP), and recall as TP/(TP+FN). The multiclass labelled ground truth masks and the predicted masks were converted to one-hot encoding labels to calculate the per-lesion TP, TN, FP, and FN.
Figure 3: Visualization of the ground truth and predictions of RICAU-Net, U-Net, and Nested U-Net on three input images. The zoomed-in images of the lesions are displayed on the right top and left bottom corners. The first row images only contain CAC in LM, the second row images contain CAC in LM and LAD, and the third row images contain CAC in LCX and RCA. In the first row, the per-lesion Dice scores of each model are LM: 100% for RICAU-Net; LM: 0% for U-Net and Nested U-Net. U-Net and Nested U-Net misclassified CAC in LM pixels as CAC in LAD. In the second and third rows, the per-lesion Dice scores of each model are LM: 74.68%, LAD: 100%, LCX: 99.74%, RCA: 97.60% for RICAU-Net; LM: 23.03%, LAD: 99.89%, LCX: 99.48%, RCA: 97.60% for U-Net; LM: 60.29%, LAD: 96.72%, LCX: 99.48%, RCA: 98.41% for Nested U-Net. Our model significantly improves the segmentation performance for sparse and smaller CAC lesions, while also enhancing the results of other major ones.
4.3 Comparison with Other Methods
The lesion-specific segmentation performance of RICAU-Net was evaluated by comparing it with six other methods: U-Net [29], SA U-Net [10], Nested U-Net [44], ResUNet++ [18], Attention U-Net [26], and VGG16 U-Net. The six methods were trained using the same datasets and implementation details as RICAU-Net, with the exception of ResUNet++. ResUNet++ used a batch size of 8. The results are shown in Tab. 1. The proposed method achieved the highest per-lesion Dice scores for all four lesions: 59.35 (LM), 94.11 (LAD), 87.03 (LCX), and 92.89 (RCA). For CAC in LM, considered as small and the scarcest lesions in our datasets, the top three methods with the highest LM Dice score were RICAU-Net, Nested U-Net, and U-Net. Among them, RICAU-Net outperformed Nested U-Net by 1.85% and U-Net by 6.33%. For precision, RICAU-Net exhibited the highest scores in two lesions of 55.14 (LM) and 86.30 (LCX). The highest precision of CAC in LAD and CAC in RCA were achieved by Nested U-Net and SA U-Net respectively, with improvements of 0.11% and 2.01% compared to RICAU-Net. However, SA U-Net achieved the lowest RCA recall of 55.79 and the lowest RCA Dice score of 69.44. This indicates that SA U-Net had a higher number of false negative pixels of CAC in RCA than false positive pixels, with the former being 9.03 times higher. For recall, RICAU-Net achieved the highest scores in two lesions of 94.23 (LAD) and 96.06 (RCA). The highest recall of CAC in LM and CAC in LCX were achieved by Nested U-Net and ResUNet++, which were 2.18% and 4.18% higher than RICAU-Net, respectively. However, ResUNet++ achieved the lowest LCX precision of 62.65 and the second lowest LCX Dice score of 74.52. This indicates that ResUNet++ had a higher number of false positive pixels of CAC in LCX compared to false negative pixels, with the former being 6.81 times higher. The segmentation performances of the RICAU-Net, U-Net, and Nested U-Net with respect to the ground truth were visualized in Fig. 3. For simplicity, only these top three methods with the highest LM Dice score were selected for visualization. Overall, RICAU-Net demonstrated superior and more consistent results compared to other methods, particularly for the rarest lesion (CAC in LM).
Table 1: Lesion-specific segmentation results of RICAU-Net and six other methods. Per-lesion Dice score, precision, and recall were compared. The other methods were reproduced by ourselves, training them using the same datasets and conditions as RICAU-Net. The segmentation performances of the background and bone were not listed in the table because the per-lesion performance of the background and bone was above 99% in most methods. Methods that exhibit per-lesion performance of the background and bone lower than 99% were marked with an asterisk ($\ast$). The highest per-lesion score in each metric is in bold.
Metrics Method LM LAD LCX RCA
@@ 4.3 Comparison with Other Methods
U-Net 53.68 93.77 81.72 88.96
SA U-Net 19.02 85.11 69.76 91.94
ResUNet++* 36.00 90.29 62.65 84.92
Precision (%) Attention U-Net 50.47 92.63 79.84 89.31
Nested U-Net 50.69 94.11 77.76 89.75
VGG16 U-Net 44.58 93.54 80.29 86.71
RICAU-Net (ours) 55.14 94.00 86.30 89.93
U-Net 52.38 93.48 86.68 93.92
SA U-Net* 46.42 70.50 66.06 55.79
ResUNet++ 58.07 88.03 91.95 92.20
Recall (%) Attention U-Net 54.20 93.37 88.23 92.54
Nested U-Net 66.44 92.74 87.84 94.09
VGG16 U-Net 65.15 90.16 85.74 87.28
RICAU-Net (ours) 64.26 94.23 87.77 96.06

### v2

@@ Abstract
The Agatston score, which is the sum of the calcification in the four main coronary arteries, has been widely used in the diagnosis of coronary artery disease (CAD). However, many studies have emphasized the importance of the vessel-specific Agatston score, as calcification in a specific vessel is significantly correlated with the occurrence of coronary heart disease (CHD). In this paper, we propose the Residual-block Inspired Coordinate Attention U-Net (RICAU-Net), which incorporates coordinate attention in two distinct manners and a customized combo loss function for lesion-specific coronary artery calcium (CAC) segmentation. This approach aims to tackle the high class-imbalance issue associated with small and sparse CAC lesions. Experimental results and the ablation study demonstrate that the proposed method outperforms the four other U-Net based methods used in medical applications, by achieving the highest per-lesion Dice scores across all four lesions.
Index Terms—  Coordinate attention, Small lesion segmentation, Coronary artery calcium segmentation
@@ 1 Introduction
The Agatston score [1], also known as coronary artery calcium (CAC) score, is widely used as a non-invasive method for prediction of coronary artery disease (CAD) by using non-contrast cardiac computed tomography (CT) [2, 3, 4]. The CAC score is calculated as the sum of calcification in the four main coronary arteries: the left main coronary artery (LM), left anterior descending artery (LAD), left circumflex artery (LCX), and right coronary artery (RCA). With the advancement of artificial intelligence in the medical field, deep learning algorithms have shown its capability of locating the calcium in the coronary arteries and automating the calculation of the CAC score with high precision [5, 6, 7]. Although the total Agatston score has been widely used for decades, many studies [5, 7, 8] emphasized the importance of the lesion-specific CAC scores. The authors in [9] concluded that the presence of CAC in RCA was significantly correlated with coronary heart disease (CHD) and cardiovascular disease (CVD). Therefore, the importance of the per-lesion Agatston score is as significant as the conventional Agatston score in predicting CHD and CVD.
To calculate the per-lesion Agatston score, the lesion-specific CAC segmentation is a mandatory process. However, there are two limitations when facilitating the lesion-specific segmentation of CAC. The first factor to consider is the size and sparsity of the lesions. CAC is located in the coronary arteries; therefore, the size of CAC is relatively smaller than that of other organs such as the lung, kidney, or liver. The second issue is the high class-imbalance problem in the datasets. CAC in LM is the least common lesion among the four main coronary arteries because LM is the shortest coronary artery [10]. Therefore, the number of CT slices where CAC in LM appeared is the smallest, and the size of the LM lesion is also small. Many studies [5, 6, 7, 8] focused on lesion-specific segmentations, but [6, 7] combined CAC in LM and LAD as one lesion due to the challenge of distinguishing CAC in LM from CAC in LAD.
@@
Herein, we propose the Residual-block Inspired Coordinate Attention U-Net (RICAU-Net), an end-to-end deep learning algorithm for lesion-specific CAC segmentation using non-contrast cardiac CT scans. We also propose a customized combo loss function to tackle the segmentation of small and sparse lesions with a high class-imbalance problem. In the reporting deep learning algorithm, coordinate attention (CA) [11] was implemented as RICA blocks in the encoder and as CA modules in the decoder of the proposed model. This implementation aimed to highlight the positional information of CAC as the locations of CAC were closely related to their surrounding vessels in CT images, which is substantial for the lesion-specific segmentation task. Through this study, we demonstrate the superior performance of the proposed algorithm and loss function compared to other U-Net based networks for lesion-specific CAC segmentation.
Fig. 1: Overview of the RICAU-Net architecture (left), and the CA module and the RICA block in RICAU-Net (right). The CA module illustrates the overall process of aggregating positional information from both horizontal and vertical directions of a feature map into channel attention. In the RICA block, the final output is obtained by performing an element-wise addition operation of the output of the two convolutional layers and the output of the skip connection.
2 Methods
@@ 2.1 Coordinate Attention Module
The CA module generates direction-aware and position-sensitive feature maps by breaking down channel attention into two 1D feature encoding processes that retain information in both horizontal and vertical directions [11]. These feature maps help the neural network accurately detect and locate regions of interest in the input image. Hence, the CA module was incorporated into the proposed architecture to emphasize the positional information of CAC and its surrounding arteries from the input images. The CA module is defined as:
@@ 2.2 Residual-block Inspired Coordinate Attention U-Net
The decoder of the RICAU-Net is similar to that of the vanilla U-Net, but it incorporates four CA modules after each upsampling process. In this way, the network is capable of retaining the positional information of the upsampled features. The RICA block was not utilized in the decoder of the RICAU-Net because it could result in the loss of essential positional information when the feature maps from the CA module pass through the PJS module for dimension matching.
@@ 2.3 Class-imbalance Problem and Loss Function
2.3.1 Class-imbalance Problem
Our non-contrast cardiac CT scan dataset shows a significant class imbalance between the negative samples (background, bone) and the positive samples (CAC in LM, LAD, LCX, RCA). For instance, Among training set, the smallest number of CAC was found in LM (1.3%), whereas the largest number of CAC was observed in RCA (7.4%). This phenomenon persisted in the validation and test sets as well. Among the validation and test sets, only 1.3% and 0.4% of images contained CAC in LM, respectively. It was more significant at the pixel level. In a $512\times 512$ pixel image, one LM lesion could be as small and sparse as 5 pixels $(>1mm^{2})$, whereas the remaining pixels would belong to the negative samples.
2.3.2 Customized Combo Loss Function
Our dataset suffers from two major problems. The first issue is high class imbalance. The second issue is the small and sparse sizes of the target lesions. Focal loss [13] was first introduced to address the class imbalance problem in the object detection field. Exponential Logarithmic Loss (ELL) [14] was introduced to address the highly unbalanced object size problem in 3D segmentation. ELL is a type of combo loss that is based on a mixture of Cross-entropy loss and Dice loss. To overcome the issues, we proposed a weighted Focal LogDice loss, which combined the weighted Focal loss and the Dice loss part of ELL with an exponential parameter. The weighted Focal LogDice loss is defined as:
@@ 2.3.2 Customized Combo Loss Function
where $\mathcal{L}_{WFLogDL}$ denotes the weighted Focal LogDice loss, $\mathcal{L}_{WFL}$ denotes weighted focal loss, $\mathcal{L}_{ELDL}$ denotes exponential logarithmic Dice loss, $w_{Focal}$ and $w_{Dice}$ denote weights of each loss function, and $w_{Focal}=0.4$ and $w_{Dice}=0.6$ were used in this study.
Fig. 2: Visualization of the segmentation performance of RICAU-Net, U-Net, and Nested U-Net on three input images. The zoomed-in lesion images and corresponding per-lesion Dice scores with highlights are shown in the images. Our model significantly improves the segmentation performance for sparse and smaller CAC lesions, while also enhancing the results of other major ones.
3 Experiments and Results
3.1 Dataset
A total of 1,108 non-contrast cardiac CT scans were used in this study, acquired using four different scanners: SIEMENS, GE Medical Systems, PHILIPS, and TOSHIBA. The calcium lesions in the coronary arteries were manually annotated by five clinically trained specialists to create lesion-specific ground truth masks. The ground truth masks contained multiclass information on background, bone, CAC in LM, LAD, LCX, and RCA. A total of 612 CT scans from SIEMENS, GE Medical Systems, and Philips scanners were used for the training and validation sets. The training set (80%) comprised 494 scans (484 subjects with CAC and 10 subjects without CAC), and the validation set (20%) comprised 118 scans (112 subjects with CAC and 6 subjects without CAC). 496 CT scans from TOSHIBA scanner were used for the test set, comprising 278 subjects with CAC and 218 subjects without CAC. All the CT scans had the same height and width dimensions ($512\times 512$ pixel sizes).
3.2 Preprocessing
All the CT scans were sliced into 2D images with a resolution of $512\times 512$ pixels. The raw Hounsfield Units (HU) of each slice were clipped to a range of [-150, 230], which represents the clinical CT windowing of the heart (window width: 380, window level: 40), followed by normalization to a range of floating-point numbers [0, 1]. In addition, random rotations of 5 and 10 degrees clockwise and counterclockwise, random center crops of $300\times 300$ and $400\times 400$ pixel resolutions, as well as the addition of Gaussian blur, Gaussian noise, and salt and pepper noise, were employed as on-the-fly data augmentation methods to augment the training data.
3.3 Implementation Details
@@
Table 1: The per-lesion segmentation results of RICAU-Net and four other methods. The per-lesion Dice scores of the background and bone were above 99% in all methods. The highest per-lesion Dice scores are in bold.
Method Per-lesion Dice Scores (%)
LM LAD LCX RCA
@@
3.4 Evaluation Metric
Dice score was used as the evaluation metric. It is one of the most commonly used evaluation metrics in semantic segmentation tasks because it measures the similarity between the ground truth and the model’s prediction.
3.5 Comparison with Other Methods
The lesion-specific segmentation performance of RICAU-Net was evaluated by comparing it with four other methods: U-Net [17], Nested U-Net [18], Attention U-Net [19], and VGG16 U-Net. The four methods were trained using the same datasets and implementation details as RICAU-Net. The results are shown in Table 1. The proposed method achieved the highest per-lesion Dice scores for all four lesions: 59.35 (LM), 94.11 (LAD), 87.03 (LCX), and 92.89 (RCA). For CAC in LM, considered as small and the scarcest lesions in our datasets, the top three methods with the highest LM Dice score were RICAU-Net, Nested U-Net, and U-Net. Among them, RICAU-Net outperformed Nested U-Net by 1.85% and U-Net by 6.33%. The segmentation performances of these methods were visualized in Fig. 2. Overall, RICAU-Net demonstrated superior and more consistent results compared to other methods for all four lesions.